OpenStudy (anonymous):

PLEASE HELP :C Solve for x: the quantity of x plus 4 over 3 = 2. x = −2 x = 2 x = 2 over 3 x = negative 10 over 3

OpenStudy (texaschic101):

(x + 4) / 3 = 2......is this your problem... OR.. x + (4/3) = 2...or this one

OpenStudy (anonymous):

First one sorry

OpenStudy (texaschic101):

(x + 4) / 3 = 2 --- multiply both sides by 3...this cancels out the 3 on the left side x + 4 = (3)(2) x + 4 = 6 -- subtract 4 from both sides x = 6 - 4 x = 2

OpenStudy (anonymous):

Thankyou :)

OpenStudy (texaschic101):

of course :) thats what I am here for